# Graphene-hBN interlayer interactions from quantum Monte Carlo

### v2

@@ I.1 Quantum Monte Carlo
AB $(s=0)$Refer to caption SP $(s=1/6)$Refer to caption Mid $(s=1/2)$Refer to caption AA $(s=2/3)$Refer to caption
@@
AB $(s=0)$Refer to caption SP $(s=1/6)$Refer to caption BA $(s=1/3)$Refer to caption AA $(s=2/3)$Refer to caption
Figure 2: Stacking registries for rigid bilayer graphene and graphene-hBN bilayer. In both systems, the lattice constants of both layers are set to the monolayer graphene lattice constant $a_{\textrm{G}}=2.46~{}\textrm{\AA}$. The SP, BA, Mid, and AA stacking types are defined by translating the top layer while fixing the bottom layer by $s=1/6,1/3,1/2$, and $2/3$, respectively in the units of $\sqrt{3}a_{\textrm{G}}$.
@@
where $E(\infty)$ is a fitting parameter, which represents the extrapolated energy, and $N$ is the number of primitive cells in a simulation cell. The statistical error bar from QMC is obtained from the bootstrapping technique.
In our graphene-hBN bilayer simulation cell, hBN is compressed to the lattice constant of graphene $a_{\textrm{G}}=2.46~{}\textrm{\AA}$. We tested that the effect of this compression is negligible by performing calculations at different lattice constants (Fig. 3 and Fig. 6). QMC energy data points are collected from configurations AB, SP, BA, and AA stacking registries (Fig. 2) with interlayer distances from 3 to $8~{}\textrm{\AA}$ (Fig. 4).
@@ I.2.1 Intralayer potential
Figure 3: The in-plane compression and stretching of the graphene-hBN bilayer from QMC, DFT-D2, DFT-D3, and the atomistic calculation using the REBO, ExTeP, and ILP-Ouyang potentials. The vertical lines denote the lattice constants of monolayer graphene $a_{\textrm{G}}=2.46~{}\textrm{\AA}$ and monolayer hBN $a_{\textrm{hBN}}=2.504~{}\textrm{\AA}$. The line fit is a Morse potential. The statistical error bars from QMC are smaller than the markers.
@@ I.2.2 Interlayer potential
where $R_{\mathrm{cut}}$ is fixed to $16~{}\mathrm{\text{\AA}}$ throughout this work.
@@ III Conclusion
IV Appendix

### v3

@@ I.1 Quantum Monte Carlo
AB (s=0)(s=0)Refer to caption SP (s=1/6)(s=1/6)Refer to caption Mid (s=1/2)(s=1/2)Refer to caption AA (s=2/3)(s=2/3)Refer to caption
@@
AB (s=0)(s=0)Refer to caption SP (s=1/6)(s=1/6)Refer to caption BA (s=1/3)(s=1/3)Refer to caption AA (s=2/3)(s=2/3)Refer to caption
[…]
@@
where $E(\infty)$ is a fitting parameter, which represents the extrapolated energy, and $N$ is the number of primitive cells in a simulation cell. For each supercell, we performed twist averaging on the aforementioned $\times 4$ k-point grid. The statistical error bar from QMC is obtained from the bootstrapping technique.
[…]
@@ I.2.1 Intralayer potential
[…]
@@ I.2.2 Interlayer potential
[…]
@@ III Conclusion
IV Acknowledgments
This project is supoprted by U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Computational Materials Sciences Award No. DE-SC0020177. An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used resources from the Argonne Leadership Computing Facility, a U.S. DOE Office of Science user facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility (project qmc-hamm). This work used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC0500OR22725.
V Appendix